Chromium metal can be plated out from an acidic solution containing CrO3 according to the following reaction:CrO3 + 6H++ 6e–  Cr + 3H2OCalculate the mass of chromium plated out by 24000 coulomb. How long will it take to plate out 1.5 gof chromium using 12.5 ampere current?

Grade:12th pass

1 Answers

Arun
3 years ago
1. charge Q = 
 
where n is number of moles of the species under consideration, is the n-factor [ simply n-factor means change in oxidation number ]
 
and F is charge required to deposit 1 mole of electron i.e., 96500C
 
here, n - factor = 6
 
Charge, Q = 24000C
 
so, 24000C = n × 6 × 96500C
 
or, n = 24000/(96500 × 6) = 0.04145mol
 
now, mole = weight in gm/atomic weight
 
so, 0.04145 = weight in gm/52g/mol
 
or, weight in gm of chromium = 0.04145 × 52 = 2.1554g
 
2. number of mole of Cr = 1.5/52
 
number of mole of electron passed = 6 × 1.5/52 [ because change in oxidation of Cr is +6 ]
 
we know, 1 mole of electron is equivalent to 96500C charges.
 
so, charge required = 6 × 1.5/52 × 96500
 
= 9 × 96500/52
 
= 16702C
 
time required = charge/current
 
= 16702/12.5
 
= 1336.16 sec.